# Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

### v1

@@ 5.1 Experiment Settings
Dataset. Three distinct datasets are utilized: the alignment dataset, harmful dataset, and fine-tuning dataset. The alignment dataset and harmful dataset are derived from the RepNoise (Rosati et al., 2024b), which extracts subsets from the BeaverTails dataset (Ji et al., 2023). Specifically, 5,000 examples are sampled for the alignment dataset, and 1,000 examples for the harmful dataset. The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K (Cobbe et al., 2021), SST2 (Socher et al., 2013), AlpacaEval (Li et al., 2023), and AGNEWS (Zhang et al., 2015), with 1,000 samples collected from each task (700 samples from AlpacaEval). To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with $(1-p)$ of benign fine-tuning data, and $p$ is set to 0.1 by default. The harmful data is also sourced from BeaverTails (Ji et al., 2023) and does not overlap with the harmful dataset.
@@
Baseline. Four methods are considered as baselines in our experiments. The SFT method is the vanilla supervised training using the alignment dataset. Vaccine (Huang et al., 2024e) applies supervised training on the alignment dataset while introducing additional perturbations. Both RepNoise (Rosati et al., 2024b) and Booster (Huang et al., 2024b) utilize the alignment and harmful datasets for supervised and adversarial training. All four baseline methods are trained exclusively during the alignment stage. Our proposed method incorporates training with the harmful dataset during the fine-tuning stage after vanilla alignment training. More details are in Appendix C.
@@
Metric. Following  (Huang et al., 2024e), two metrics are used to evaluate the model’s performance.
@@
  • HS (Harmful Score): It reflects the frequency with which the model generates harmful content when handling malicious instructions. Harmful Score is determined by a moderation model, provided by  (Ji et al., 2023), which assesses whether the model’s output is harmful in response to a given harmful instruction. And a sample of 1,000 instructions is drawn from the BeaverTails (Ji et al., 2023) test set to compute this metric.
@@
  • FA (Finetuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks. Samples are sampled from the test sets of GSM8K, SST2, AlpacaEval, and AGNEWS, with 1,000, 872, 1,000, and 105 samples, respectively, used to compute this metric. More details are in Appendix B.
@@
Implementation Details. For efficient training, the approach follows the methodology (Huang et al., 2024e), utilizing LoRA (Hu et al., 2021) with a rank of 32 and an alpha value of 4. And the optimizer is AdamW (Loshchilov et al., 2017). During the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the training is performed for 20 epochs. For the fine-tuning stage, the learning rate is set to $52e-$ 5, with a batch size for 10 and a training epoch for 20. These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K (Cobbe et al., 2021) and the default model being Llama2-7B (Touvron et al., 2023) following (Rosati et al., 2024b; Huang et al., 2024b). To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024), are included in the evaluation.
@@
SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.68 59.97
Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.20 57.36
RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.00 59.71
Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.05 60.36
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.88 59.66
5.2 Main Results
In this sub-section, we conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
Harmful Ratio. Fine-tuning datasets with different harmful ratios are employed, specifically 0 (clean), 0.05, 0.1, 0.15, 0.2, and 0.25. The results are presented in Table 1, Panacea achieves the lowest harmful score across different harmful ratios while maintaining competitive fine-tuning performance (ranked as the second-best on average), indicating that the expected adaptive perturbation is obtained, and the analysis is in Sec 5.3. Compared to SFT method, it reduces the harmful score by an average of 21.5% and improves fine-tuning accuracy by 0.3%. Furthermore, as the harmful ratio increases, Panacea consistently maintains a lower harmful score compared to other methods. By mitigating the impact of harmful loss, the model achieves the best fine-tuning performance. However, since Panacea is designed to weaken harmful loss, its fine-tuning performance on clean data (without explicit harmful loss) is slightly reduced, while Panacea still achieves the lowest harmful score on this benign fine-tuing.
@@ 5.3 Perturbation Analysis
Our post-fine-tuning perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and in some cases, it even enhances model performance as reducing harmful behaviors.
Table 4: Ablation study on different fine-tuning tasks. “w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
@@
Table 5: Ablation study on different LLMs. ”w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
@@
Figure 5: Parameter weights of different LLMs. The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen (purple) have larger weights in the middle and later layers.
Visualization. Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 5. The result in Figure 5 reveals the layers of different LLMs that are most critical for model’s safety. Specifically, in Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting that the earlier layers are more critical for the model’s safety. This observation aligns with previous research (Du et al., 2024a; Rosati et al., 2024b; Liu et al., 2024a; Yi et al., 2024b). Additionally, we observe that in Gemma2-9B, the middle and final layers hold greater importance for safety, while in Qwen2-7B, the safety importance gradually increases across layers, reaching its peak in the final layers. This may indicate that the model implements stricter safety measures in the output layer. Therefore, our experiments also suggest that these models may require targeted defense mechanisms for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
@@ 5.4 Statistical Analysis
Harmful Score. As shown in the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process. However, the harmful score increase for Panacea is generally smaller. In the final few hundred steps, the defense of SFT is significantly degraded, causing a sharp rise in harmful scores, while the harmful score of Panacea remains stable. This is because, at this stage, the perturbation optimized by Panacea enhances the model’s ability to mitigate harmful behaviors.
Harmful Training Loss. As shown in the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT. Although both harmful training losses decrease, Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end. With the relatively high loss, Panacea’s harmful score shows a significant improvement. Finally, we observe an interesting phenomenon that the harmful training loss and harmful score of Panacea follow the same trend in the last few hundred steps. We attribute this to the near-successful optimization of the adaptive perturbation at this stage. See more analysis of harmful loss in Appendix B.3.
Harmful Testing Loss. As shown in the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss. This is because the data used during training and testing come from the same distribution, and the methods applied during training are transferable to unseen data.
5.5 System Evaluation
Table 6 presents a comparison of the clock time and GPU memory usage during training for different methods.
Table 6: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. ”Align.” and ”Fine.” suggest alignment and fine-tuning.
Methods Clock Time (Hour) GPU Memory (GB)
Align. Fine. Sum Align. Fine. Max
SFT 0.58 0.17 0.75 34.90 32.90 34.90
Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
Booster 1.87 0.17 2.04 35.98 32.90 35.98
Panacea 0.58 0.42 1.00 34.90 32.86 34.90
@@ 5.6 Hyper-parameter Analysis
Regularizer Intensity $\lambda$. In Eq. 1, by adjusting $\lambda$, we ensure that the harmful loss is increase after adding the perturbation. Table 8 shows the impact of regularizer intensity $\lambda$ on Panacea. It can be observed that $\lambda$ is linearly related to both the harmful score and fine-tuning accuracy. As $\lambda$ increases, the harmful score decreases more, but fine-tuning performance also drops. This is in expectation, because with a larger $\lambda$, we put more weights in maximize the harmful loss after adding perturbation, and therefore resulting in a lower harmful score. On the other hand, we do observe that when $\lambda$ is too large, the fine-tune accuracy will degrade. Therefore, a validation dataset is required to select a proper $\lambda$ based on the need of the deployment.
Table 7: Impact of perturbation intensity $\rho$ over Panacea. As $\rho$ increases, the harmful score decreases. However, when $\rho$ continues to increase, the harmful score decreases further, but the fine-tuning accuracy also begins to decrease.
@@
Table 8: Impact of regularizer’s intensity $\lambda$ over Panacea. As $\lambda$ increases, the harmful score and the fine-tuning accuracy decrease.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to address the harmful fine-tuning and make LLMs helpful and harmless. We acknowledge that the phenomena or issues identified in this paper may pose potential risks. Disclaimer: this paper contains red-teaming data (from open dataset) and model-generated content that can be offensive in nature.

### v2

@@ 5.1 Experiment Settings
Dataset. Three distinct datasets are utilized: the alignment dataset, harmful dataset, and fine-tuning dataset. The alignment dataset and harmful dataset are derived from the RepNoise Rosati et al. (2024b), which extracts subsets from the BeaverTails dataset Ji et al. (2023). Specifically, 5,000 examples are sampled for the alignment dataset, and 1,000 examples for the harmful dataset. The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K Cobbe et al. (2021), SST2 Socher et al. (2013), AlpacaEval Li et al. (2023), and AGNEWS Zhang et al. (2015), with 1,000 samples collected from each task (700 samples from AlpacaEval). To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with (1p)(1-p) of benign fine-tuning data, and $p$ is set to 0.1 by default. The harmful data is also sourced from BeaverTails Ji et al. (2023) and does not overlap with the harmful dataset.
@@
Baseline. Four methods are considered as baselines in our experiments. The SFT method is the vanilla supervised training using the alignment dataset. Vaccine Huang et al. (2024e) applies supervised training on the alignment dataset while introducing additional perturbations. Both RepNoise Rosati et al. (2024b) and Booster Huang et al. (2024b) utilize the alignment and harmful datasets for supervised and adversarial training. All four baseline methods are trained exclusively during the alignment stage. Antidote Huang et al. (2024a)is a post-fine-tuning stage baseline that utilizes a harmful dataset after the fine-tuning stage. Our proposed method incorporates training with the harmful dataset during the fine-tuning stage after vanilla alignment training. More details are in Appendix E.
@@
Metric. Following  Huang et al. (2024e), two metrics are used to evaluate the model’s performance.
@@
  • HS (Harmful Score): It reflects the frequency with which the model generates harmful content when handling malicious instructions. Harmful Score is determined by a moderation model, provided by  Ji et al. (2023), which assesses whether the model’s output is harmful in response to a given harmful instruction. And a sample of 1,000 instructions is drawn from the BeaverTails Ji et al. (2023) test set to compute this metric.
@@
  • FA (Fine-tuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks. Samples are sampled from the test sets of GSM8K, SST2, AlpacaEval, and AGNEWS, with 1,000, 872, 1,000, and 105 samples, respectively, used to compute this metric. Details are in Appendeix D.1
@@
Implementation Details. For efficient training, the approach follows the methodology Huang et al. (2024e), utilizing LoRA Hu et al. (2021) with a rank of 32 and an alpha value of 4. And the optimizer is AdamW Loshchilov et al. (2017). During the alignment stage, the learning rate is set to 5e45e-4, the batch size is 10, and the training is performed for 20 epochs. For the fine-tuning stage, the learning rate is set to 2e52e-5, with a batch size for 10 and a training epoch for 20. These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K Cobbe et al. (2021) and the default model being Llama2-7B Touvron et al. (2023) following Rosati et al. (2024b); Huang et al. (2024b). To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B Team et al. (2024) and Qwen2-7B Yang et al. (2024a), are included in the evaluation.
5.2 Main Results
We conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
Harmful Ratio. Fine-tuning datasets with different harmful ratios are employed, specifically 0 (Clean), 0.05, 0.1, 0.15, 0.2. The results are presented in Table 1, Panacea achieves the lowest harmful score across different harmful ratios while maintaining competitive fine-tuning performance (ranked as the second-best on average), indicating that the expected adaptive perturbation is obtained, and the analysis is in Sec 5.3. Compared to SFT method, it reduces the harmful score by an average of 21.2% and improves fine-tuning accuracy by 0.4%. Furthermore, as the harmful ratio increases, Panacea consistently maintains a lower harmful score compared to other methods. By mitigating the impact of harmful loss, the model achieves the best fine-tuning performance. However, since Panacea is designed to weaken harmful loss, its fine-tuning performance on clean data (without explicit harmful loss) is slightly reduced, while Panacea still achieves the lowest harmful score on this benign fine-tuing.
Table 1: Performance comparison of different harm ratio.
Method Harmful Score($\downarrow$) Fine-tuning Accuracy($\uparrow$)
Clean 0.05 0.1 0.15 0.2 Avg. Clean 0.05 0.1 0.15 0.2 Avg.
SFT 5.2 27.5 45.8 56.2 67.0 40.3 16.4 16.0 16.2 15.4 15.2 15.8
Vaccine 2.3 15.6 25.4 40.3 55.2 27.8 14.2 13.9 13.5 13.0 13.6 13.6
RepNoise 2.7 22.2 32.0 42.9 54.0 30.8 15.7 16.1 15.8 14.9 14.0 15.3
Booster 4.6 21.0 42.3 60.7 69.3 39.6 17.8 17.8 17.6 17.1 16.2 17.3
Antidote 2.3 14.0 27.2 32.7 35.3 22.3 17.9 16.2 15.3 16.3 16.2 16.3
Panacea 1.8 9.9 20.1 29.1 34.8 19.1 15.0 16.3 16.7 17.0 16.2 16.2
@@ 5.2 Main Results
SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.7 59.9
Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.2 57.3
RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.0 59.7
Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.0 60.3
Antidote 27.2 15.3 43.5 93.58 19.4 33.01 47.5 84.9 34.4 56.6
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.9 59.6
@@
Aligned LLMs. Since our method operates at the post-fine-tuning stage, it can be directly applied to already safety-aligned large language models (LLMs), such as Llama2-7B-Chat, Gemma2-9B-It, and Qwen2-7B-Instruct. In contrast, methods like Vaccine must be applied during the alignment stage and are therefore inapplicable to pre-aligned LLMs. We compare our approach with two other post-fine-tuning methods: Safe LoRA Hsu et al. (2024) and Antidote Huang et al. (2024a).
As shown in Table 4, Panacea reduces the average harmful score by about 20% compared to the SFT baseline, while incurring only a 0.2% drop in accuracy. Moreover, compared to other post-fine-tuning stage methods, our method consistently achieves greater reductions in harmful scores. On the Gemma model, Panacea achieves nearly a 15% lower harmful score than the best competing method, demonstrating the effectiveness of the optimized perturbation introduced by our approach.
Table 4: Performance comparison on aligned LLMs.
Llama2-7B-Chat Gemma2-9B-It Qwen2-7B-Instruct Average
Method HS FA HS FA HS FA HS FA
SFT 47.2 20.0 54.4 77.0 28.9 67.2 43.5 54.7
Safe LoRA 46.8 20.5 56.8 76.3 29.9 67.5 44.5 54.8
Antidote 37.3 20.4 31.0 76.9 23.3 59.3 30.5 52.2
Panacea 35.7 17.2 15.6 80.3 19.8 66.1 23.7 54.5
Evaluation Benchmark. To better validate the effectiveness of Panacea, we conduct additional evaluations on both Sorry-Bench Xie et al. (2025) and the AdvBench Zou et al. (2023) under different harmful ratios using LLaMA-2-7B. We also include the ConstrainSFT method that is proposed in Qi et al. (2024a) Section 4.1. The evaluation results are presented in Table 5. For Sorry-Bench, Fulfillment Rate (FR) is used as the metric (lower is better ↓). As the results show, Panacea consistently achieves the best performance across all three settings, demonstrating its effectiveness and generalizability. In particular, on AdvBench, the harmful score remains as low as 10.58% even under the most extreme setting.
Table 5: Evaluations on AdvBench and Sorry-Bench under diffusion harmful ratios.
AdvBench HS (ratio=0.05) HS (ratio=0.1) HS (ratio=0.15) HS (ratio=0.2)
SFT 7.50 17.69 37.50 48.65
Vaccine 25.19 49.62 59.23 71.35
RepNoise 3.46 8.46 20.38 40.00
ConstrainSFT 4.04 12.12 20.58 34.81
Panacea 0.00 1.54 5.19 10.58
Sorry-Bench FR (ratio=0.05) FR (ratio=0.1) FR (ratio=0.15) FR (ratio=0.2)
SFT 45.23 57.50 65.00 70.23
Vaccine 34.32 49.77 53.63 65.45
RepNoise 60.91 66.59 75.23 81.14
ConstrainSFT 45.23 59.55 61.36 66.36
Panacea 33.41 42.05 49.55 54.32
Harmful Data from Different Sources. We conducted the experiment that harmful data is from different sources. Specifically, the harmful data used during the defense phase remains from BeaverTails Ji et al. (2023), while the harmful data used for fine-tuning in the attack phase is replaced with data from LLM-LAT Sheshadri et al. (2024). And the harmful score is evaluated using test set from AdvBench Zou et al. (2023). The experimental results are shown in Table 6. As shown, Panacea significantly reduces the harmful score compared to other methods, even when the harmful data come from different sources. This result further demonstrates the effectiveness of our method.
Table 6: Performance comparison under different harmful data sources.
Method HS (ratio=0.05) HS (ratio=0.1) FA (ratio=0.05) FA (ratio=0.1)
SFT 74.62 84.81 16.5 15.9
Vaccine 48.65 73.65 15.2 13.5
RepNoise 61.73 81.54 15.4 14.7
ConstrainSFT 60.00 86.35 15.0 15.5
Panacea 11.73 41.73 17.1 17.1
@@ 5.3 Perturbation Analysis
Table 7: Ablation study on different fine-tuning tasks. “w/o" denotes that the adaptive perturbation obtained is not applied after the training.
@@
Table 8: Ablation study on different LLMs. "w/o" denotes that the adaptive perturbation obtained is not applied after the training.
@@
Our perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and it even enhances model performance as reducing harmful behaviors.
Layer-wise safety property. Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 8. We sum the absolute values of the parameters at each layer, to compute the magnitude of changes by perturbation to the original model.
Figure 5: Parameter weights of different LLMs. The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen2-7B (purple) have larger weights in the middle and later layers.
The result in Figure 5 reveals different layers exhibit different affinity towards safety task. In Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting the earlier layers are more critical for the model’s safety. This observation aligns with multiple previous research Du et al. (2024a); Rosati et al. (2024b); Liu et al. (2024a); Yi et al. (2024b); Li et al. (2024a); Zhao et al. (2025b) (See Appendix C for detailed discussions). Additionally, we observe that in Gemma2-9B, the middle and final layers hold greater importance for safety, while in Qwen2-7B, the safety importance gradually increases across layers, reaching its peak in the final layers. This may indicate that the model implements stricter safety measures in the output layer. Therefore, our experiments also suggest these models may require targeted defense for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
@@ 5.4 Statistical Analysis
Harmful Score. In the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process. However, the harmful score increase for Panacea is generally smaller. In the final few hundred steps, the defense of SFT is significantly degraded, causing a sharp rise in harmful scores, while the harmful score of Panacea remains stable. This is because, at this stage, the perturbation optimized by Panacea enhances the model’s ability to mitigate harmful behaviors.
Harmful Training Loss. In the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT. Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end. With the high loss, Panacea’s harmful score shows a significant improvement. Finally, we observe an interesting phenomenon that the harmful training loss and harmful score of Panacea follow the same trend in the last few hundred steps. We attribute this to the near-successful optimization of the adaptive perturbation at this stage. See more analysis of harmful loss in Appendix D.3.
Harmful Testing Loss. In the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss. This is because the data used during training and testing come from the same distribution, and the methods applied during training are transferable to unseen data.
@@ 5.5 Hyper-parameter Analysis
Regularizer Intensity $\lambda$. In Eq. 1, by adjusting $\lambda$, we ensure that the harmful loss is increase after adding the perturbation. Table 10 shows the impact of regularizer intensity $\lambda$ on Panacea. It can be observed that $\lambda$ is linearly related to both the harmful score and fine-tuning accuracy. As $\lambda$ increases, the harmful score decreases more, but fine-tuning performance also drops. This is in expectation, because with a larger $\lambda$, we put more weights in maximize the harmful loss after adding perturbation, and therefore resulting in a lower harmful score. Besides, we observe that when $\lambda$ is too large, the fine-tune accuracy will degrade.
Table 9: Impact of perturbation intensity $\rho$.
@@
Table 10: Impact of regularizer’s intensity $\lambda$.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
Li Shen is supported by National Key R&D Projects (NO. 2024YFC3307100), NSFC Grant (No. 62576364), Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO. JCYJ20241202124430041), CCF-DiDi GAIA Collaborative Research Funds (NO. CCF-DiDi GAIA 202419 and CCF-DiDi GAIA 202519)